find the perimeter of the square. (one side=2x) (other side=x+25)
Answer by stanbon(75887)

find the perimeter of the square. (one side=2x) (other side=x+25)
---
The sides of a square are all equal.
Equation
2x = x+25
x = 25
----
Perimeter = 4*side = 4x = 4*25 = 100
=====================================
Cheers,
Stan H.